you earned a 75% on a test you answered 21 questions correctley how many questions were on test

Found 2 solutions by jim_thompson5910, checkley79:
Answer by jim_thompson5910(35256)

21/x = 75/100

21*100 = 75x

2100 = 75x

75x = 2100

x = 2100/75

x = 28

So there were 28 questions total on the test.

Answer by checkley79(3341)

21/.75=28 ans.